# How Can We Effectively Expand the Vocabulary of LLMs with 0.01GB of Target Language Text?

### v2

@@ 5 Training Strategy
Continual pre-training on $\mathcal{D}$ is an integral part of vocabulary expansion to improve the alignment of the newly initialized embeddings. To this end, we explore: (i) the training procedure, (ii) the objective function, and (iii) the input sequence length.
Procedure.
We use by default low-rank adaptation (LoRAHu et al. (2022) applied to all linear layers, following previous work on vocabulary expansion in high-resource settings Cui et al. (2023); Balachandran (2023); Abbasi et al. (2023); Choi et al. (2024).
We also consider the two-stage tuning process (2-stage) Cui et al. (2023), where only the embeddings and LM head are first updated, followed by tuning the LoRA modules. This process can help minimize the risk of overfitting to the initial embedding state, which might be suboptimal Downey et al. (2023). Alternatively, we train only the top and bottom two layers (2x2 LS) following Remy et al. (2024). This approach aims to calibrate only the parts closely related to the encoding and decoding of the target language, minimizing changes to the base model. Note that none of these approaches have ever been compared to the standard LoRA or against each other. We tune the embeddings and LM head in each case.
Objective.
We first evaluate clm, which has been used by previous work in high-resource settings (Cui et al., 2023; Balachandran, 2023; Larcher et al., 2023, inter alia), as a default training objective. We also consider a multi-token prediction (mtp) objective Gloeckle et al. (2024), where the model must predict multiple consecutive tokens at each timestep rather than a single token. mtp has been proposed for pre-training and exhibited performance gains over clm. However, it has yet to be explored for continual pre-training for cross-lingual transfer. We experiment with one additional LM head (i.e. predicting two consecutive tokens at a time) and initialize the additional weights with those from the original LM head. [^5] We find in our preliminary analysis that random initialization does not work well (Appendix E.3).
Sequence Length.
Finally, we shorten the default input sequence length from 2,048 to 512, thereby increasing the number of training batches. We hypothesize that this is critical in low-resource settings, as a small number of model updates could be prone to underfitting.
@@ 6.1 Source Models
We use Llama2 (7B) Touvron et al. (2023), an English-centric model, with its $\mathcal{T}_{\text{s}}$ based on byte-fallback Byte Pair Encoding (BPE) Sennrich et al. (2016) and $|\mathcal{V}_{\text{s}}|$ set to 32K in the experiments. We also use Llama3 (8B) Dubey et al. (2024) and Gemma2 (9B) Riviere et al. (2024) for consistency in our analysis. These models have a far larger vocabulary size than Llama2, i.e. 128K and 256K, respectively.
@@ 6.2 Target Languages and Data
We experiment with a typologically diverse set of ten languages with various scripts. This includes German (Indo-European) and Swahili (Niger–Congo) for the Latin script, and Arabic (Afroasiatic), Burmese (Sino-Tibetan), Greek (Indo-European), Hindi (Indo-European), Japanese (Japonic), Sinhala (Indo-European), Telugu (Dravidian), and Thai (Kra–Dai) for non-Latin scripts. We select these languages because of the availability of downstream task datasets with the same task formulation across languages.
To simulate a realistic low-resource adaptation scenario, we follow Yong et al. (2023) in using 30K sentences per language. These sentences are randomly sampled from their language-specific subcorpus of CC-100 Conneau et al. (2020).
@@
LAPT .18.00 360.2 .11.00 240.3 .20.00 210.1 .21.00 250.3 .21.00 380.2 .08.00 230.2 .07.00 274.1 .12.00 300.1 .09.00 290.1 .12.00 230.2
+Speedup Random .07.00 320.3 .05.00 170.3 .28.00 230.3 .13.00 200.1 .09.00 350.2 .18.00 220.2 .07.00 290.1 .13.00 290.1 .06.00 260.2 .07.00 190.2
@@
Table 1: Mean performance on mt and sum tasks over five runs in low-resource settings (30K sentences) using Llama2 as source. The baselines without vocabulary expansion are in gray. Darker blue and red shades indicate higher positive and negative relative performance change over Source per language and task, respectively.
6.3 Baselines
We use the following two methods as our baselines: (1) the source model $\mathcal{M}_{\text{s}}$ off-the-shelf without any adaptation (Source); (2) we continue pre-training Source on $\mathcal{D}$ by keeping $\mathcal{V}_{\text{s}}$ (LAPT). These do not offer any inference speedups.
6.4 Tasks
We evaluate all models using two tasks per target language in a zero-shot setting: (1) English-to-target machine translation (mt) using FLORES-200 NLLB Team et al. (2022) and (2) summarization (sum) including German MLSUM Scialom et al. (2020), GreekSUM Evdaimon et al. (2023), and XL-Sum Hasan et al. (2021) for the rest.666sum performance is not directly comparable between different languages as the data does not match across languages. Following Ahia et al. (2023), we use 500 random samples. The prompt templates can be found in Appendix B.
6.5 Evaluation Metrics
Task Performance.
We use chrF Popović (2015) for mt and ROUGE-L Lin (2004) for sum. We report average performance over five different runs throughout the experiments and analysis.
Perplexity.
We report perplexity on the 100K language-specific held-out sentences sampled from the CC-100 subcorpus as an auxiliary metric for evaluating model performance.
Inference Efficiency.
We measure inference efficiency as the number of tokens generated per second Hong et al. (2024).
6.6 Implementation Details
We set $|\mathcal{V}_{\text{aux}}|$ to 50K across languages and the number of new target tokens $k$ to 100 by default. We investigate the effect of varying $k$ in §7.4. For LAPT, we train each model for two epochs with a batch size of 8, a maximum learning rate of 1e-4, and a sequence length of 2,048. We set a LoRA rank to 8 following previous work Cui et al. (2023); Abbasi et al. (2023); Lin et al. (2024).
7 Results and Analysis
7.1 Target Vocabulary Initialization
We analyze the effect of different target vocabulary initialization methods (§4) on task performance and inference efficiency using Llama2 as source.
@@ 7.1 Target Vocabulary Initialization
Table 3: Inference efficiency measured by the number of tokens generated per second (T/S). The first row of Source corresponds to its respective T/S for each language and task. Bold and underlined indicate the best and second-best speedups across tasks and models with the same base model for each language.
Task Performance and Perplexity.
Table 1 shows the performance of all methods, while Table 2 shows the corresponding perplexities on the held-out language-specific dataset. We first find that models adapted with Mean and Align show comparable or better performance than Source in the majority of the cases. Specifically, they outperform or perform the same as Source in 13 and 14 out of 20 cases, respectively. This result aligns with the perplexity results in Table 2, where Align tends to exhibit the lowest perplexity across languages, followed by Mean. Further, some performance gaps in the rest of the cases are within only a few points ($\leq$ 3). In contrast, Merge shows worse performance than Align and Mean, outperforming or performing the same as Source in 11 cases. We speculate that Merge might be less informative than these two methods, as it does not rely on surface information for initialization, resulting in lower performance.
Random performs the same as or better than Source in 12 cases, while FOCUS achieves this in 11 cases—the worst of the adapted models along with Merge. This is also evident in the perplexity results, where FOCUS shows the worst perplexity in six languages (i.e. my, de, hi, si, te, and th). These results suggest that sophisticated initialization approaches do not always work well, and the popular Random is not the best choice in low-resource settings. Thus, we conclude that models adapted with Mean and Align are more likely to perform competitively to Source.
Next, we analyze the results by language and task. We see that the adapted models, especially Mean and Align, generally perform competitively with Source within a few points of degradation across all languages and substantially outperform it in Greek, Japanese and Swahili mt, with up to 17 points of improvement in Japanese. However, they often do not perform competitively in Burmese, German, Greek, and Hindi sum. We hypothesize that sum might require more training data than mt to perform well as generating long text (up to 128 generated tokens and using long context) requires models to have strong generative capabilities in a target language. We later address this challenge in §7.2 by showing that these performance gaps can be drastically narrowed using our alternative training strategies, further contributing to the competitiveness of vocabulary expansion approaches to Source.
Inference Efficiency.

### v3

@@ 5 Training Strategy
Continual pre-training on $\mathcal{D}$ (i.e. LAPT) is an integral part of vocabulary expansion to improve the alignment of the newly initialized embeddings. To this end, we explore: (i) the training procedure, (ii) the objective function, and (iii) the input sequence length.
Training Procedure
  • LoRA: We use by default low-rank adaptation (LoRAHu et al. (2022) applied to all linear layers, following previous work on vocabulary expansion in high-resource settings Cui, Yang, and Yao (2023); Balachandran (2023); Abbasi et al. (2023); Choi et al. (2024).
  • 2-stage: We also consider the two-stage tuning process (2-stage) Cui, Yang, and Yao (2023), where only the embeddings and language modeling head are first updated, followed by tuning the LoRA modules. This process can help minimize the risk of overfitting to the initial embedding state, which might be suboptimal Downey et al. (2023).
  • 2x2 LS: We train only the top and bottom two layers (2x2 LS) of the model, following Remy et al. (2024). This calibrates only the parts closely related to the encoding and decoding of the target language Wendler et al. (2024); Tang et al. (2024), minimizing changes to the source model.
Note that none of these approaches have ever been compared to the standard LoRA or against each other. We tune the embeddings and language modeling head in each case.
Objective Function
  • clm: We first evaluate a causal language modeling objective (clm), which has been used by previous work in high-resource settings (Cui, Yang, and Yao, 2023; Balachandran, 2023; Larcher et al., 2023, inter alia), as a default training objective.
  • mtp: We also consider a multi-token prediction (mtp) objective Gloeckle et al. (2024), where the model must predict multiple consecutive tokens at each timestep rather than a single token. mtp has been proposed for pre-training and exhibited performance gains over clm. However, it has yet to be explored for continual pre-training for cross-lingual transfer. We experiment with one additional language modeling head (i.e. predicting two consecutive tokens at a time) and initialize the additional weights with those from the original language modeling head. [^8] We find in our preliminary analysis that random initialization does not work well.
Input Sequence Length
We shorten the default input sequence length from 2,048 to 512, thereby increasing the number of training batches. We hypothesize that this is critical in low-resource settings, as a small number of model updates could be prone to underfitting.
@@ 6.1 Source Models
We use Llama2 7B Touvron et al. (2023), an English-centric non-instruction-tuned model, with its $\mathcal{T}_{\text{s}}$ based on byte-fallback Byte Pair Encoding (BPE) Sennrich, Haddow, and Birch (2016) and $|\mathcal{V}_{\text{s}}|$ set to 32K in the experiments. We also use Llama3 8B Dubey et al. (2024) and Gemma2 9B Riviere et al. (2024) for consistency in our analysis. These models have a far larger vocabulary size than Llama2, i.e. 128K and 256K, respectively. Note that Llama2 and Llama3 have untied embedding and language modeling head weights, while the original configuration of Gemma2 includes weight tying.
@@ 6.2 Target Languages and Data
We experiment with a typologically diverse set of ten languages with various scripts. This includes German (Indo-European) and Swahili (Niger–Congo) for the Latin script, and Arabic (Afroasiatic), Burmese (Sino-Tibetan), Greek (Indo-European), Hindi (Indo-European), Japanese (Japonic), Sinhala (Indo-European), Telugu (Dravidian), and Thai (Kra–Dai) for non-Latin scripts. We select these languages because of the availability of downstream task datasets with the same task formulation across languages, with a particular focus on generation tasks.
To simulate a realistic low-resource adaptation scenario, we follow Yong et al. (2023) in using 30K sentences per language. With $|\mathcal{D}|$ set to 30K, this equates to up to approximately 5M tokens. These sentences are randomly sampled from their language-specific subcorpus of CC-100 Conneau et al. (2020). Note that previous work on vocabulary expansion for generative LLMs Tejaswi, Gupta, and Choi (2024); Mundra et al. (2024) uses at least 200M tokens and 2.5B tokens respectively, which is at least 40 times larger than our training budget. [^9] For additional context, the BabyLM challenge Warstadt et al. (2023); Hu et al. (2024), where participants must pre-train a model from scratch under low-resource settings, utilizes training budgets of up to 100M words, motivated by the observation that children are exposed to fewer than 100M words by 13 years of age. While the BabyLM challenge is a challenging task, our setting, with access to less than 5M tokens, represents an even more extreme low-resource scenario, posing a distinct challenge for effective target language adaptation.
6.3 Baselines
We use the following two methods as our baselines:
  1. 1.
    Source: We use the off-the-shelf source base (i.e. non-instruction-tuned) model $\mathcal{M}_{\text{s}}$ without any adaptation, following Tejaswi, Gupta, and Choi (2024). This provides a crucial reference point, allowing us to quantify the inherent performance in the target language before language-specific tuning, and thereby clearly measure specific gains from vocabulary expansion and LAPT.
  2. 2.
    CPT-Only: We continue pre-training (CPT) Source on $\mathcal{D}$, retaining its original vocabulary $\mathcal{V}_{\text{s}}$. This differs from vocabulary expansion approaches, which expand the original vocabulary $\mathcal{V}_{\text{s}}$ to include new terms $\mathcal{V}_{\text{t}}=\mathcal{V}_{\text{s}}\cup\mathcal{V}_{\text{new}}$ as mentioned in §3.
It is important to emphasize that these baselines do not offer any inference speedups.
6.4 Evaluation
Tasks
We use both generation and classification target language tasks to evaluate each approach. For generation tasks, we use (1) English-to-target machine translation (mt) using FLORES-200 (NLLB Team et al., 2022) and (2) summarization (sum) including German MLSUM Scialom et al. (2020), GreekSUM Evdaimon et al. (2024), and XL-Sum Hasan et al. (2021) for the rest. Note that sum performance is not directly comparable between different languages as the data does not match across languages. For classification tasks, we employ multiple-choice reading comprehension (mc) using Belebele (Bandarkar et al., 2024) and Global MMLU (gmmlu(Singh et al., 2025) as a general knowledge and reasoning benchmark. Note that gmmlu does not support Burmese and Thai.
Number of Samples
Following Ahia et al. (2023), we use 500 random samples for generation tasks (sum and mt). mc and gmmlu use their full test sets for evaluation. Specifically, mc has 900 samples per language subset, while gmmlu has 14K samples.
Prompt Templates
For sum prompt templates, we translate the English templates from Ahia et al. (2023) using a machine translation API, as in Yong et al. (2023). For mt, we create an English template and then translate it into each target language. For mc and gmmlu, we follow the default template provided by Hugging Face LightEval Habib et al. (2023). The complete prompt templates are listed in Table 12 in the Appendix.
6.5 Evaluation Metrics
Task Performance
We use accuracy for mc and gmmlu, chrF Popović (2015) for mt, and ROUGE-L Lin (2004) for sum. In the analysis (§7.3), we also use BLEURT Sellam, Das, and Parikh (2020) as an auxiliary metric for sum.
We report average zero-shot performance across five different runs for the generation tasks, namely sum and mt. For the classification tasks, we report single-run three-shot performance for mc and five-shot performance for gmmlu as these tasks are deterministically evaluated with temperature set to zero.
Perplexity
We report perplexity on 100K language-specific held-out CC-100 sentences as an auxiliary metric for evaluating model performance.
Inference Efficiency
We measure inference efficiency as the number of tokens generated per second Hong, Lee, and Cho (2024).
6.6 Implementation Details
Hyperparameters
We set $|\mathcal{V}_{\text{aux}}|$ to 50K across languages and the number of new target tokens $|\mathcal{V}_{\text{new}}|$ to 100 by default. We investigate the effect of varying $|\mathcal{V}_{\text{new}}|$ in §7.4. For LAPT, we train each model for two epochs with a batch size of 8, a maximum learning rate of 1e-4, and a sequence length of 2,048. We set a LoRA rank to 8 following previous work Cui, Yang, and Yao (2023); Abbasi et al. (2023); Lin et al. (2024). Table 13 in the Appendix details the hyperparameter configurations employed during both the training and inference phases.
To make a fair comparison, we do not conduct any parameter tuning and use the same ones across all approaches. For sum, we truncate an article whenever it exceeds the maximum prompt length of 4,096 to avoid the CUDA out-of-memory error.
Evaluation Metrics Computation
To compute ROUGE-L, we split sentences with an mT5 Xue et al. (2021) tokenizer as preprocessing following Maynez, Agrawal, and Gehrmann (2023) and subsequently call rouge_scorer [^10] https://github.com/csebuetnlp/xl-sum/tree/master/multilingual_rouge_scoring to compute the metric. We use BLEURT-20 to compute BLEURT. [^11] https://github.com/google-research/bleurt
Libraries and Hardware
For Llama2 and Gemma2, we train tokenizers using SentencePiece Kudo and Richardson (2018) and convert them into the Hugging Face Tokenizers Moi and Patry (2023) format. For Llama3, we train tokenizers using Hugging Face Tokenizers. We implement our models using PyTorch Paszke et al. (2019), Hugging Face Transformers Wolf et al. (2020) and PEFT Mangrulkar et al. (2022). We preprocess datasets with Hugging Face Datasets Lhoest et al. (2021). For evaluation, we use Hugging Face LightEval Habib et al. (2023). We use either four NVIDIA V100 (32GB) or a single A100 (80GB) for LAPT. Evaluation utilizes a single NVIDIA V100 (32GB) or A100 (80GB). Each analysis utilizes a consistent hardware configuration to ensure accurate measurement of inference efficiency.
7 Results and Analysis
7.1 Target Parameter Initialization
We analyze the effect of different target parameter initialization methods (§4) on task performance and inference efficiency using Llama2 as source.
7.1.1 Task Performance and Perplexity
Table 1 show the performance of all methods on generation tasks, while Table 2 shows the corresponding perplexities on the held-out language-specific dataset.
Performance on Generation Tasks
Models initialized with Mean and Align generally exhibit performance comparable to or better than Source. Specifically, they outperform (by >> 2 points) or match (within 2 points) Source in 16 and 15 out of 20 cases, respectively. Mean shows a positive gain over Source in 10 cases, while Align achieves this in 12 cases. These results aligns with their perplexity scores (Table 2), where Align generally yields the lowest perplexity across languages, closely followed by Mean. While Merge also matches or surpasses Source performance in 16 cases (with 10 positive gains), it often shows the largest perplexities among the three heuristic-based initialization methods. We speculate that Merge might be less informative than these two methods, as it does not rely on surface information for initialization, resulting in slightly larger perplexity.
Table 1: Mean performance on generation tasks (mt and sum) over five runs in low-resource settings (30K sentences) using Llama2 as source. Green shades indicate positive performance change over Source per language and task, respectively.
@@ 7.1.1 Task Performance and Perplexity
CPT-only .18.00 360.2 .11.00 240.3 .20.00 210.1 .21.00 250.3 .21.00 380.2 .08.00 230.2 .07.00 230.1 .12.00 300.1 .09.00 290.1 .12.00 230.2
+Speedup
Random .07.00 320.3 .05.00 170.3 .28.00 230.3 .13.00 200.1 .09.00 350.2 .18.00 220.2 .07.00 290.1 .13.00 290.1 .06.00 260.2 .07.00 190.2
@@
The baseline Random and FOCUS models perform similarly (within 2 points) to or better than Source in 14 cases, with 10 and 7 cases showing positive gains, respectively. These results place them among the least effective approaches. FOCUS, in particular, shows the fewest cases with positive gains and the worst perplexity in six languages (i.e. Burmese, German, Hindi, Sinhala, Telugu, and Thai). These findings suggest that sophisticated initialization methods do not always guarantee superior performance, possibly due to underfitting of an auxiliary embedding model. Further, the popular Random approach is not always optimal in low-resource scenarios. Thus, we conclude that models initialized with Mean and Align are more likely to perform competitively with Source on generation tasks.
Analyzing performance by language and task, Mean and Align models show positive gains over Source in six and eight out of ten languages for mt, respectively. In cases where gains are not observed, any performance degradation is negligible (within 2 points). Notably, they substantially outperform Source in Greek, Japanese, and Swahili mt, with improvements ranging from 8 to 17 points.
However, for sum, these adapted models often do not yield a performance gain over Source in the majority of the languages, except for Japanese, Sinhala, Swahili, and Telugu. In particular, Burmese and Greek show a substantial performance drop of up to 9 and 8 points, respectively. We hypothesize that sum may necessitate more training data than mt because generating longer text (up to 128 tokens, using long context) requires models to have strong generative capabilities in the target language. We later address this challenge in §7.2 by showing that these performance gaps can be drastically narrowed using our alternative training strategies, further contributing to the competitiveness of vocabulary expansion approaches to Source.
Turning to CPT-only (i.e. continual pre-training without vocabulary expansion), it demonstrates strong performance, matching or improving upon Source in 17 out of 20 cases, a higher count by at least one case than achieved by the models with the heuristic-based initialization. A direct comparison against the best-performing Align [^12] It is best-performing because it generally achieves the lowest perplexities across languages and is the most likely among different initialization methods to provide a positive gain over Source. further confirms this. CPT-only outperforms Align in 10 out of 20 cases, whereas Align only wins in four. These results demonstrate the overall superiority of CPT-only over vocabulary expansion approaches. This trend aligns with previous work on CVA Downey et al. (2023); Yamaguchi, Villavicencio, and Aletras (2024) suggesting that CPT-only can often perform better than vocabulary expansion approaches in low-resource settings, possibly due to its reliance on robust and well-aligned original embeddings.
Table 3: Mean performance on classification tasks (mc and gmmlu) in low-resource settings (30K sentences) using Llama2 as source. Green shades indicate positive performance change over Source per language and task, respectively. Note that gmmlu does not support Burmese and Thai.
mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu
Source .29 .29 .26 - .43 .39 .27 .28 .25 .28 .40 .33 .24 .27 .31 .28 .28 .27 .29 -
CPT-only .30 .29 .22 - .42 .38 .29 .29 .28 .28 .39 .34 .23 .27 .33 .27 .24 .27 .28 -
+Speedup
Random .28 .25 .22 - .42 .38 .22 .27 .27 .26 .35 .30 .22 .27 .29 .25 .28 .26 .28 -
FOCUS .29 .26 .28 - .41 .38 .24 .26 .28 .26 .36 .31 .29 .26 .28 .25 .24 .27 .28 -
Mean .28 .25 .23 - .42 .38 .30 .26 .27 .26 .38 .29 .28 .26 .29 .25 .23 .27 .30 -
Merge .29 .25 .25 - .43 .38 .27 .28 .26 .28 .37 .29 .27 .26 .28 .25 .28 .26 .29 -
Align .28 .25 .26 - .40 .37 .31 .26 .27 .27 .37 .28 .27 .27 .29 .25 .22 .27 .31 -
Performance on Classification Tasks
Table 3 presents the performance of all methods on classification tasks. A distinct trend emerges compared to generation tasks: adapted models rarely yield a positive performance gain over Source. While all vocabulary expansion approaches exhibit competitive or better performance than Source in at least 12 cases, their positive gains are limited to a maximum of four cases (22%) (Mean and Align). This is far fewer than the maximum of 12 cases (60%) observed on the generation tasks.
Notably, CPT-only performs on par with or outperforms Source in 16 out of 18 cases. Nonetheless, this contrasts with its strong performance on generation tasks: it shows positive gains in only 6 out of 18 cases (33%), compared to 65% on generation tasks. These results suggest that while continual pre-training (i.e. LAPT) can offer benefits, its impact on discriminative tasks in low-resource settings is limited and differs from its effect on generation tasks.
We hypothesize these differences arise because our adaptation process (LAPT) primarily optimizes a causal language modeling objective on unlabeled target language data $\mathcal{D}$. While this objective is highly beneficial for generative tasks requiring fluency and extended text production, it may not directly translate to immediate gains on classification tasks. Classification tasks, in contrast, often heavily depend on the inherent semantic and factual knowledge of a model, frequently requiring only a single token generation for prediction. Therefore, while previous work on vocabulary expansion with LAPT Tejaswi, Gupta, and Choi (2024); Cui, Yang, and Yao (2023); Balachandran (2023); Choi et al. (2024); Yamaguchi et al. (2025) often reports performance improvements in both generation and classification tasks in resource-rich settings, our findings suggest that under extremely low-resource settings (i.e. 30K sentences per language, approximately up to 5M tokens), LAPT does not inherently guarantee an improvement in these discriminative capabilities.
Table 4: Inference efficiency for: (a) generation tasks (mt and sum) over five runs; and (b) classification tasks (mc and gmmlu). We measure inference efficiency by the number of tokens generated per second (T/S). The Source row shows both raw T/S and baseline 1.00x speedup per language and task. Bold and underlined indicate the best and second-best speedups across tasks and models with the same base model for each language.
(a) Generation tasks
@@
(b) Classification tasks
mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu
19.4 29.9 11.1 - 36.6 34.3 14.0 23.4 16.9 24.1 33.4 34.0 10.9 15.2 35.4 33.1 11.1 12.8 16.7 -
Source 1.00x 1.00x 1.00x - 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x -
CPT-only 1.17x 0.98x 1.24x - 1.04x 0.97x 1.22x 0.92x 1.17x 0.99x 1.03x 0.93x 1.27x 1.10x 1.04x 1.02x 1.02x 1.11x 1.20x -
Random 1.71x 1.12x 2.12x - 1.03x 1.01x 1.68x 1.29x 1.71x 1.25x 1.02x 0.99x 2.10x 1.92x 1.05x 1.01x 1.84x 2.10x 1.42x -
FOCUS 1.71x 1.12x 2.13x - 0.99x 1.01x 1.68x 1.29x 1.71x 1.23x 1.02x 0.99x 2.11x 1.92x 1.05x 1.00x 1.84x 2.11x 1.42x -
Mean 1.71x 1.10x 2.13x - 1.02x 1.01x 1.68x 1.29x 1.71x 1.25x 1.03x 1.00x 2.10x 1.92x 1.05x 1.03x 1.84x 2.11x 1.42x -
Merge 1.71x 1.12x 2.12x - 1.02x 0.99x 1.68x 1.29x 1.71x 1.25x 1.02x 0.99x 2.10x 1.93x 1.06x 1.01x 1.84x 2.10x 1.42x -
Align 1.71x 1.11x 2.12x - 1.03x 1.02x 1.68x 1.29x 1.71x 1.25x 1.03x 0.98x 2.10x 1.92x 1.06x 1.03x 1.84x 2.10x 1.42x -
7.1.2 Inference Efficiency